if f(x+1)=(x-2)(x-1) then find f(x)

Grade:Upto college level

3 Answers

Prudhvi teja
13 years ago

dear mohd

let x+1 =y

f(y) = (y-3)(y-2)

 

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian

rama rao
13 years ago

replace x with x-1 in f(x+1)=(x-1)(x-2) then u will get f(x)=(x-2)(x-3)

Shashank Shekhar
13 years ago

Replace x by x-1,

thus,f(x)=(x-3)(x-2).